OpenStudy (anonymous):

what percentage of 46400 is 169?

OpenStudy (anonymous):

try to divide 169/46400

OpenStudy (anonymous):

let it be x percentage (x/100) *46400 = 169 x=169/464 = 0.36422 %

OpenStudy (anonymous):

Al you do is divide the two